# Jiawei Zhang    Shuang Yang    Bo Li

### v1

@@ 5 Ablation Study
Different Number of Locations for Noise Insertion.
Varying the number of locations for noise insertion affects the pattern of successful attacks, as demonstrated in Table 4 for the InjecAgent dataset. While Joint optimization generally delivers similar performance when using different number of locations, but if success is counted whenever at least one setting (ranging from 1 to 4 locations) succeeds, the ASR for direct harm reaches 98%, and 94% for data stealing, which shows that there is actually little overlap among the successful cases. Additionally, as shown in Table 5, using more locations for noise insertion on the AgentHarm dataset demonstrates enhanced attack efficiency.
Different Optimization Modes.
In scenarios involving malicious observation, such as InjecAgent as shown in Table 4, Sequential Optimization proves more effective than Joint Optimization. Conversely, in scenarios of malicious instruction like AgentHarm—where the model may refuse the request—Joint Optimization significantly outperforms Sequential Optimization, as detailed in Table 5.
Attack Efficiency.
As indicated in Figure 3, we calculate the average number of optimization iterations required for a successful attack with UDora on a subset where attacks are consistently successful across different numbers of locations. It is observed that whether using Sequential or Joint Optimization, no more than 30 iterations are necessary to achieve a successful attack on an LLM agent, showcasing extreme efficiency of UDora. Notably, on AgentHarm, UDora with Joint Optimization requires fewer than 10 iterations on average to conduct a successful attack, which demonstrates the importance of using models’ own reasoning style to attack themselves.
@@ 6 Red-Teaming on Real World Agents
Web Search Agent.
Nowadays, the performance gap between closed and open-source models is narrowing. Powerful open-source models like DeepSeek R1 (Guo et al., 2025) have been increasingly deployed by many companies, such as Perplexity, Microsoft, and Dell, in their applications. This widespread deployment exposes more vulnerabilities to transferability attacks. Take Perplexity AI’s Web Search Agent Sonar (Perplexity, 2024) as an example, a possible adversarial attack in this scenario could be like “how to do money laundering.” In other words, an adversary could compel the agent to gather related information from the internet and provide corresponding references. We find that by leveraging the optimized adversarial strings developed for AgentHarm on the Llama 3.1 model can easily be adapted to attack the Sonar agent, detailed examples are provided in Section C.1.
AI Email Agent.
AutoGen recently introduced an email agent (AutoGen, 2024) that can autonomously generate email replies via the Gmail tool from Composio, based on the GPT-4o model from OpenAI. By exploiting only the log probabilities returned by GPT-4o, and applying greedy string optimization based on positional scoring without gradient backpropagation using UDora, we successfully attacked this email agent. The attack involved sending emails containing optimized adversarial strings, which led the agent to disclose the content of all recent emails or personal profile information, which is highly sensitive. This occurred even with the agent being prompted not to reveal any private information. Detailed attack logs and the optimization procedure are deferred to Section C.2.
All these two cases underscores the importance of red-teaming in the real-world deployment of LLM agents.
@@ 7 Discussion
In this work, we introduce UDora, a unified red-teaming framework designed to attack LLM agents by exploiting their own reasoning styles with high attack efficiency. Despite its effectiveness, there are areas for potential improvement in future work. For instance, our framework requires access to the token probability distribution for a successful attack; however, in some cases, only the pure text response is available, which can undermine our attack strategy. Moreover, in our current methodology, we consistently use the same noise for insertion across different positions. Employing a more diverse range of noises on different positions could potentially mislead the LLM agent more effectively, enhancing the overall efficacy of the attacks.
@@ Acknowledgment
This work is partially supported by the National Science Foundation under grant No. 2046726, NSF AI Institute ACTION No. IIS-2229876, DARPA GARD, the National Aeronautics and Space Administration (NASA) under grant No. 80NSSC20M0229, ARL Grant W911NF-23-2-0137, the Alfred P. Sloan Fellowship, the Meta research award, the AI Safety Fund, and the eBay research award.
Impact Statements
@@
Appendix A Experiment Details
Appendix B Example Processes on Optimizing Adversarial Strings Across Multiple Datasets
B.1 Optimization Process on InjecAgent
@@
B.2 Optimization Process on WebShop
@@
B.3 Optimization Process on AgentHarm
@@
Appendix C Red-Teaming on Real World Agents
C.1 Attack on Perplexity AI based on transferbility
@@
C.2 Attack on the AI Email Agent from AutoGen

### v2

@@ 5 Ablation Study
Different Number of Locations for Noise Insertion. Varying the number of locations for noise insertion affects the pattern of successful attacks, as demonstrated in Table 4 for the InjecAgent dataset. While Joint optimization generally delivers similar performance when using different number of locations, but if success is counted whenever at least one setting (ranging from 1 to 4 locations) succeeds, the ASR for direct harm reaches 98%, and 94% for data stealing, which shows that there is actually little overlap among the successful cases. Additionally, as shown in Table 5, using more locations for noise insertion on the AgentHarm dataset demonstrates enhanced attack efficiency.
Different Optimization Modes. In scenarios involving malicious observation, such as InjecAgent as shown in Table 4, Sequential Optimization proves more effective than Joint Optimization. Conversely, in scenarios of malicious instruction like AgentHarm—where the model may refuse the request—Joint Optimization significantly outperforms Sequential Optimization, as detailed in Table 5.
Attack Efficiency. As indicated in Figure 3, we calculate the average number of optimization iterations required for a successful attack with UDora on a subset where attacks are consistently successful across different numbers of locations. It is observed that whether using Sequential or Joint Optimization, no more than 30 iterations are necessary to achieve a successful attack on an LLM agent, showcasing extreme efficiency of UDora. Notably, on AgentHarm, UDora with Joint Optimization requires fewer than 10 iterations on average to conduct a successful attack, which demonstrates the importance of using models’ own reasoning style to attack themselves. To further evaluate computational efficiency, we find that the main per-iteration cost stems from optimizing the adversarial string (similar to the GCG attack (Zou et al., 2023)) rather than the response collection (query). For example, on the AgentHarm dataset with LLaMA 3.1, the average times per iteration are 3.48 seconds for query and 11.22 seconds for optimization, while on InjecAgent, these are 5.57 and 5.97 seconds, respectively. With the Ministral model, the corresponding times are 2.49 and 6.04 seconds on AgentHarm, and 6.86 and 8.18 seconds on InjecAgent. In practice, updating the reasoning less frequently (e.g., every 10 steps) or generating only the first 100 reasoning tokens per query can further reduce computational overhead.
@@ 6 Red-Teaming on Real World Agents
Web Search Agent. Nowadays, the performance gap between closed and open-source models is narrowing. Powerful open-source models like DeepSeek R1 (Guo et al., 2025) have been increasingly deployed by many companies, such as Perplexity, Microsoft, and Dell, in their applications. This widespread deployment exposes more vulnerabilities to transferability attacks. Take Perplexity AI’s Web Search Agent Sonar (Perplexity, 2024) as an example, a possible adversarial attack in this scenario could be like “how to do money laundering.” In other words, an adversary could compel the agent to gather related information from the internet and provide corresponding references. We find that by leveraging the optimized adversarial strings developed for AgentHarm on the Llama 3.1 model can easily be adapted to attack the Sonar agent, detailed examples are provided in Section B.1.
AI Email Agent. AutoGen recently introduced an email agent (AutoGen, 2024) that can autonomously generate email replies via the Gmail tool from Composio, based on the GPT-4o model from OpenAI. By exploiting only the log probabilities returned by GPT-4o, and applying greedy string optimization based on positional scoring without gradient backpropagation using UDora, we successfully attacked this email agent. The attack involved sending emails containing optimized adversarial strings, which led the agent to disclose the content of all recent emails or personal profile information, which is highly sensitive. This occurred even with the agent being prompted not to reveal any private information. Detailed attack logs and the optimization procedure are deferred to Section B.2.
@@ 7 Discussion
Limitations and Future Directions. In this work, we introduce UDora, a unified red-teaming framework designed to attack LLM agents by exploiting their own reasoning styles with high attack efficiency. Despite its effectiveness, there are areas for potential improvement in future work. For instance, our framework requires access to the token probability distribution for a successful attack; however, in some cases, only the pure text response is available, which can undermine our attack strategy. Moreover, in our current methodology, we consistently use the same noise for insertion across different positions. Employing a more diverse range of noises on different positions could potentially mislead the LLM agent more effectively, enhancing the overall efficacy of the attacks.
Potential Defenses. To mitigate potential misuse of UDora, one practical strategy is to restrict access to the agent’s internal reasoning, for example by providing only condensed or sanitized summaries of reasoning steps (as in OpenAI’s o1 reasoning model). This reduces the attacker’s ability to target reasoning traces for noise insertion. Another direction is the use of guardrail frameworks for LLM agents (Xiang et al., 2024; Luo et al., 2025), which monitor and block unsafe or abnormal behaviors during reasoning. Adversarial training (Madry et al., 2018) can also enhance robustness by exposing models to adversarially generated examples during training, though this may come with a trade-off in generalization or utility. Additionally, detection mechanisms—such as using a separate LLM to validate whether the agent’s reasoning aligns with the original instruction or environment—can help flag adversarial manipulations, though reliably distinguishing adversarial from benign reasoning remains a challenging open problem.
@@ Acknowledgment
This work is partially supported by the National Science Foundation under grant No. 1910100, No. 2046726, NSF AI Institute ACTION No. IIS-2229876, DARPA TIAMAT No. 80321, the National Aeronautics and Space Administration (NASA) under grant No. 80NSSC20M0229, ARL Grant W911NF-23-2-0137, Alfred P. Sloan Fellowship, the research grant from eBay, AI Safety Fund, Virtue AI, and Schmidt Science.
Impact Statement
@@
Appendix A Example Processes on Optimizing Adversarial Strings Across Multiple Datasets
A.1 Optimization Process on InjecAgent
@@
A.2 Optimization Process on WebShop
Examples on the attack process on WebShop on both optimization modes are shown in Figure 6 and Figure 7, where it seems that the adversarial string ultimately leads the LLM agent to hallucinate and select the wrong item (Zhang et al., 2024a). The number of location for noise insertion is also consistently set to $3$ here.
@@ A.2 Optimization Process on WebShop
A.3 Optimization Process on AgentHarm
@@
Appendix B Red-Teaming on Real World Agents
B.1 Attack on Perplexity AI based on transferbility
@@
B.2 Attack on the AI Email Agent from AutoGen